This is how many classes students at Bronx College take this semester with the probabilities

courses 1 2 3 4 5
probability (distribution) of taking that many courses .1 .1 .6 .1 .1

The variance for this distribution is approximately
Answer #1

The probability distribution is,

Courses ( X ) Probability ( P(X) ) X * P(X) X2 * P(X)
1 0.1 0.1 0.1
2 0.1 0.2 0.4
3 0.6 1.8 5.4
4 0.1 0.4 1.6
5 0.1 0.5 2.5
X * P(X) = 3 sum X^2 * P(X) = 10

The expected value of X is,

E(X)= sum X* P(X) = 3

The variance of X is,

Var(X) 〉 X) * P(X)-E(X)12 = 10-3]2 = 10-9 = 1

The variance for given distribution is approximately 1